# Sparse Autoencoders Reveal Universal Feature Spaces Across Large Language Models

### v1

@@ 6 Conclusion
In this study, we address the underexplored domain of feature universality in sparse autoencoders (SAEs) trained on LLMs with multiple layers and across multiple model pairs. To achieve these insights, we developed a novel methodology to pair features with high mean activation correlations, and then assess their global similarity of the resulting subspaces using SVCCA and RSA. Our findings reveal a high degree of similarity in SAE feature spaces across various models. Furthermore, our research reveals that subspaces of features associated with specific semantic concepts, such as calendar or people tokens, demonstrate remarkably high similarity across different models. This suggests that certain semantic feature subspaces are universally encoded across varied LLM architectures. Our work enhances understanding of how LLMs represent information at a fundamental level, opening new avenues for further research into model interpretability and the transferability of learned features.
@@
We perform analysis on available pre-trained SAEs for similar LLMs that use the same tokenizer, and preferrably on SAEs with the same number of features to match them. However, there are a limited number of pre-trained SAEs on large models. Our study is also limited by the inherent challenges in interpreting and comparing high-dimensional feature spaces. The methods we used to measure similarity, while robust, may not capture all nuances of feature interactions and representations within the SAEs. Additionally, our analysis is constrained by the computational resources available, which may have limited the scale and depth of our comparisons. The generalizability of our findings to SAEs trained on more diverse datasets or specialized domain-specific LLMs remains an open question. Furthermore, the static nature of our analysis does not account for potential temporal dynamics in feature representation that might occur during the training process of both LLMs and SAEs.
Future Work
We performed similar analysis on the LLMs that the SAEs in this paper were trained on, and found that they have some similarity, but much lower than that of the SAEs, with no patterns showing that the middle layers were the most similar. Additionally, the LLMs did not have corresponding weights for the "residual stream", which is an addition of activations from the MLP and attention head layers, so we compared the MLPs of LLMs. Unlike the SAE dimensions, the MLP dimensions of the LLMs had different dimensions. Future work could study the comparison of LLM similarity by feature weights to SAEs. Additional experiments could include using ground-truth feature experiments for LLMs that partially share ground-truth features to see if SAEs are modeling the same features (Sharkey et al., 2022), and performing functional similarity experiments (Bricken et al., 2023).
@@
Michael conceived the study and designed the experiments with help from Fazl. Michael performed the experiments with help from Austin and led the writing of the manuscript. Philp, Ashkan and David provided advice throughout the project.
Fazl advised this work and helped significantly with the write-up of the paper.
@@ Acknowledgments
We thank the Torr Vision Group at the University of Oxford for providing compute and hosting Michael Lan during the internship.
@@
We thank Clement Neo, Luke Marks, Minseon Kim, Lovis Heindrich, Tingchen Fu, Constantin Venhoff, Hazel Kim, Will Lin, Alasdair Paren, Christian Schroeder de Witt, Kiho Park, and Max Klabunde for providing helpful feedback. Clement Neo helped with the design of Figure 1.
Reproducibility
We include the code and instructions for reproducing our main results in the Supplementary Materials.
Appendix A Illustrated Steps of Comparison Method
@@
Appendix B Noise of Many-to-1 mappings
@@
Appendix C Noise of Non-Concept Features
We define "non-concept features" as features which are not modeling specific concepts that can be mapped well across models. Their highest activations are on new lines, spaces, punctuation, and EOS tokens. As such, they may introduce noise when computing similarity scores, as their removal appears to greatly improve similarity scores. We hypothesize that one reason they introduce noise is that these could be tokens that aren’t specific to concepts, but multiple concepts; thus, they are mixing concept spaces up when matching is done.
@@ Appendix C Noise of Non-Concept Features
  • <bos>
  • <|endoftext|>
@@
Appendix D Concept Categories List and Further Analysis
@@
Appendix E Additional Results

### v2

@@ 6 Conclusion
In this study, we address the underexplored domain of feature universality in sparse autoencoders (SAEs) trained on LLMs with multiple layers and across multiple model pairs. To achieve these insights, we developed a novel methodology to pair features with high mean activation correlations, and then assess their global similarity of the resulting subspaces using SVCCA and RSA. Our findings reveal a high degree of similarity in SAE feature subspaces across various models. Furthermore, our research reveals that subspaces of features associated with specific semantic concepts, such as calendar or people tokens, demonstrate remarkably high similarity across different models. This suggests that certain semantic feature subspaces are universally encoded across varied LLM architectures. However, our results also show that there are parts of SAE feature spaces which do not match, and may suggest that either certain features are not captured universally, or that SAEs have limitations in capturing universal features. We believe that future work can benefit from these discoveries to find better ways on improving methods to capture feature universality. Our work enhances understanding of how LLMs represent information at a fundamental level, opening up new avenues for further research into model interpretability and the transferability of learned features.
@@
We perform analysis only on a sample of SAEs for similar LLMs that use the same tokenizer, and focus on SAEs with the same or similar number of features. As we pair featuers by activation correlation, we do not perform analysis on models that use different tokenizers. Although, for some examples, we obtain good results when pairing featuers only by semantic similarity, these results were not consistent when applied to other examples, perhaps due to the lower pairing quality, as the representational similarity metrics depend on the pairing choice. Future work can analyze models whose representations do not depend as much on the tokenizer choice (team et al., 2024).
Our study is also limited by the inherent challenges in interpreting and comparing high-dimensional feature spaces. The methods we used to measure similarity, while robust, may not capture all nuances of feature interactions and representations within the SAEs. Additionally, our analysis is constrained by the computational resources available, which may have limited the scale and depth of our comparisons. The generalizability of our findings to SAEs trained on more diverse datasets or specialized domain-specific LLMs remains an open question. Furthermore, the static nature of our analysis does not account for potential temporal dynamics in feature representation that might occur during the training process of both LLMs and SAEs. Lastly, functional similarity experiments were not examined.
Impact Statement
Our study investigates feature universality in large language models (LLMs) using sparse autoencoders (SAEs) to identify common feature spaces across models. This research contributes to safety research by providing insights into how different LLMs encode safety-related concepts, which has several ethical and societal implications. By identifying universal feature spaces, our approach may aid in the detection of harmful or deceptive model behaviors across different architectures. Understanding how models encode knowledge could lead to improved transparency and control mechanisms, reducing risks associated with goal misgeneralization and deceptive alignment in AI systems. Additionally, identifying shared feature structures can inform bias mitigation strategies, as universal features could either amplify or help detect biases present across multiple models.
@@
Michael conceived the study and designed the experiments with help from Fazl. Michael performed the experiments with help from Austin and led the writing of the manuscript. Philp, Ashkan and David provided advice throughout the project. Fazl advised this work and helped significantly with the write-up of the paper.
@@ Acknowledgments
We thank the Torr Vision Group at the University of Oxford for providing compute and hosting Michael Lan during the internship. We thank Clement Neo, Luke Marks, Minseon Kim, Lovis Heindrich, Tingchen Fu, Constantin Venhoff, Hazel Kim, Will Lin, Alasdair Paren, Christian Schroeder de Witt, Kiho Park, and Max Klabunde for providing helpful feedback. Clement Neo helped with the design of Figure 1.
Appendix A Within Model Feature Space Similarity Results
To examine SAE feature space similarities when trained on the same model, we train multiple SAEs on a one-layer Tinystories model.
SAE Training Parameters. We train two SAEs at different seeds for a TinyStories-1L-21M model (Eldan, 2023). Both SAEs are trained at 30k steps with expansion factor of 16, for a 16384 total features, on a TinyStories dataset, and using ReLU as the activation function. We use a batch size of 4096 tokens and train with an Adam optimizer, applying a constant learning rate of $1\times 10^{-5}$. The learning rate follows a warm-up-free schedule with a decay period spanning 20% of the total training steps. L1 sparsity regularization is applied with a coefficient of 5, and its effect gradually increases over the first 5% of training steps.
We obtain SAE activations using 500 samples from the TinyStories dataset with a max sequence length of 128 (for a total of 64000 tokens).
ManyA-to-1B. This experiment allows multiple features from SAE $A$ to map onto a single feature from SAE $B$. Before filtering non-concept features, this comparison achieves an SVCCA score of 0.73, retaining 9930 feature pairs. After filtering for non-concept features (for this study, we only filter for periods and padding tokens), this pair achieves an SVCCA score of 0.94. Randomly paired achieves a score of 0.004 for 10 runs. We found that filtering for 1-1 did not change the score much.
In contrast, for 10 random runs where we randomly choose 7 tokens that are not part of the non-concept tokens, the filtering of these randomly chosen tokens does not change the unfiltered SVCCA score of 0.73.
1A-to-ManyB. This experiment allows multiple features from SAE $B$ to map onto a single feature from SAE $A$. Before filtering nonconcept features, this comparison achieves an SVCCA score of 0.8. After filtering for nonconcept features, this pair achieves an SVCCA score of 0.94, retaining 13073 feature pairs. We found that filtering for 1-1 did not change the score much. The results for "randomly chosen tokens filtering" and "randomly shuffled pairing" match those for the ManyA-to-1B experiment.
Cosine Similarity. Given that these SAEs were trained on the same model, we also compare use cosine similarity instead of activation correlation to pair the features. We obtain an average cosine similarity of 0.9. When using cosine similarity to pair the SAE features, we obtain an SVCCA score of 0.92. This occurs whether we use 1A-to-ManyB or ManyA-to-1B. Every feature finds a unique pairing, despite neither matrix containing exact matches for their features with one another.
Different SAE Widths. We find similar results for SAEs trained using an expansion factor of 8, for 8192 total features. For ManyA-to-1B, before filtering, an SVCCA score of 0.84 is achieved, and after filtering, an SVCCA score of 0.94 is achieved. For 1A-to-ManyB, before filtering, an SVCCA score of 0.87 is achieved, and after filtering, an SVCCA score of 0.94 is achieved.
Appendix B Cross Model Llama Results
We compare Llama 3-8B and Llama 3.1-8B at each model’s Layer 25 residual stream, a later layer in both models. The Llama 3 SAE has a width of 65536 features and uses Gated ReLU (Rajamanoharan et al., 2024a), while the Llama 3.1 SAE has a width of 32768 features and uses JumpReLU (Rajamanoharan et al., 2024b). We obtain activations using 100 samples with a max sequence length of 100, for a total of 10000 tokens. After filtering, this comparison achieves an SVCCA score of 0.3, comparable to the score of later layers in both Gemma and Pythia.
Appendix C Universal Vs True Features
Our definition of universal features is based on Claim 3 of "Zoom In: An Introduction to Circuits", which defines universality as studying the extent to which analogous features and circuits form across models and tasks (Olah et al., 2020). Previous work has investigated similar claims by showing evidence for highly correlated neurons (Yosinski et al., 2014; Gurnee et al., 2024) and similar feature space representations at hidden layers (Kornblith et al., 2019; Klabunde et al., 2023). Using these definitions of universality, our results find both highly correlated features and similar feature representations in SAEs.
On the other hand, “true features” (Bricken et al., 2023) can be interpreted as atomic linear directions, such that activations are a linear combination of "true feature" directions (Till, 2023), though this term has not been formally defined yet. This definition for features imposes stricter requirements which do not just rely on high correlation and similar representations, but define "true features" as linear directions that model the same concepts, assuming they exist in ground truth. Previous work has shown that SAEs may capture ground truth features Sharkey et al. (2022), while recent research has provided evidence that SAE features are not atomic (Bussmann et al., 2024), which suggests SAEs may not be sufficient to capture "true features" if they exist.
Chughtai et al. (2023a) define universality in terms of weak and strong universality. Strong universality claims that models trained in similar ways will develop the same features and circuits, whereas weak universality argues that fundamental principles, such as similar functionalities, are learned across models, but their exact implementations may vary considerably. From this perspective, strong universality would correspond more with "true features", while our paper would align closer to measuring weak universality in terms of similar feature space correlations. However, we also define analogous feature universality as a property between strong and weak universality which hypothesizes that while overall representations may greatly differ, there may be transformations that can be applied to a subset of these representations to better reveal their similarity.
The Importance of Analogous Features. While features across models may not capture the same ground truth feature representations, having analogous feature spaces could suggest that there may be transformations or mappings between the representation spaces of models which reveal their similarity. As such, this has implications for transfer learning and model stitching in which representations, including cross-layer circuits, from one model may be transferred to another (possibly larger) model through a mapping model. This may allow for functionality transfer. These studies may also have implications for furthering studies in neuroscience, as past research has been conducted on measuring the extent of alignment across biological and artificial neural networks (Goldstein et al., 2024), and previous work has also found high-low feature detectors in both vision models (Schubert et al., 2021) and mouse brains (Ding et al., 2023).
Currently, there are no standardized definitions for analogous features, though it may be possible to develop them. For instance, one possible way to better formalize a definition of analogous features would be along the lines of using a criteria akin to $(f\circ g)(x)=(h\circ f)(y)$ , given that $f$ is a mapping, while $g$ and $h$ are relations, such as distances, between points $x\in X$ and $y\in Y$ for spaces $X$ and $Y$.
This paper seeks to measure the extent of universality in terms of "analogous features in representational space", and unlike Huh et al. (2024), does not claim that models converge towards finding universal features. This follows previous papers which also measured the extent to which different neural networks learn the same representation (Wang et al., 2018a). We provide evidence that shows there are non-negligible signals of universality captured by subsets of SAE feature spaces across models, but also claim that many feature space subsets in SAEs are dissimilar in terms of their representations and relations in feature space. Still, since this may be due to the limitations of SAEs in capturing features (Engels et al., 2024b), these results do not conclude that in these dissimilar areas, models do not share universal features, and further work must be done in this area to draw stronger conclusions.
Appendix D Illustrated Steps of Comparison Method
@@
Appendix E Noise of Many-to-1 mappings
@@
Appendix F Noise of Non-Concept Feature Pairings
We define "non-concept features" as features which are not modeling specific concepts that can be mapped well across models. Their highest activations are on new lines, spaces, punctuation, and EOS tokens. As such, they may introduce noise when computing similarity scores, as their removal appears to greatly improve similarity scores. We hypothesize that one reason they introduce noise is that these could be tokens that aren’t specific to concepts, but multiple concepts. More specifically, since tokens are context dependent to a model, as information propagates through layers, the activations at a token position can aggregate information about that position’s neighbors. Therefore, it is possible that these features are not "noisy" features, but are capturing concepts which are not represented well purely by activation correlation based on highest activating positions, and thus they are not well matched via activation correlation.
@@ Appendix F Noise of Non-Concept Feature Pairings
A second hypothesis is that there are not diverse enough tokens to allow features to fire on what they should fire on. As such, providing more data to obtain correlations may allow these features to activate on more concepts, and thus be matched more accurately.
@@
  • respective padding token of tokenizer ( <bos>, <|endoftext|>, etc.)
@@
Appendix G Concept Categories List and Further Analysis
@@
Appendix H Additional Results for Pythia and Gemma